# Error Correction Code Transformer: From Non-Unified to Unified

### v1

@@ 3 Proposed Unified Error Correction Code Transformer
3.1 Unified Attention
The vanilla self-attention module can be viewed as establishing correlations among different input features within a data sample while neglecting the intrinsic relationships between different samples. This limitation hinders its ability to learn the distinctiveness of features between various code types in channel decoding tasks. Furthermore, research in [27] suggests that the output of the softmax function is probabilistically sparse, indicating that only a few weights dominate. This implies that an $N\times N$ self-attention matrix is not strictly necessary.
Based on the aforementioned considerations, we have developed a novel, low-complexity Unified Attention Module (UAM) to tackle channel decoding tasks involving multiple code types, lengths, and rates. We introduced learnable memory components, $\textbf{{A}}_{l}$ (attention) and $\textbf{{V}}_{l}$ (values), which play the role of learning features from the entire training dataset. The mathematical expression is given by
$UniAttn\left({{\textbf{{A}}_{l}},{\textbf{{V}}_{l}}}\right)=Softmax\left({{\textbf{{A}}_{l}}}\right)\times\left({{\textbf{{V}}_{l}^{T}}\textbf{{X}}}\right)$ (11)
where $\textbf{{X}}\in\mathbb{R}^{N\times d_{k}}$ is the input of attention layer, $\textbf{{A}}_{l},\textbf{{V}}_{l}\in\mathbb{R}^{N\times d_{l}}$ are trainable parameters, and $d_{l}$ is an adjustable hyperparameter.
In order to gain a better understanding of the proposed UAM, we rewrite equation (2) as follows
$Attention\left(\textbf{{W}}^{\textit{Q}},\textbf{{W}}^{\textit{K}},\textbf{{W}}^{\textit{V}}\right)=Softmax\left(\frac{\left(\textbf{{XW}}^{\textit{Q}}\right)\times{\left(\textbf{{XW}}^{\textit{K}}\right)}^{T}}{\sqrt{d_{k}}}\right)\times\left(\textbf{{XW}}^{\textit{V}}\right)$ (12)
where $\textbf{{W}}^{\textit{Q}},\textbf{{W}}^{\textit{K}},\textbf{{W}}^{\textit{V}}\in\mathbb{R}^{d_{k}\times d_{k}}$ are the projection matrices.
@@ 2 Trainable Parameters: $\textbf{{A}}_{l}$, $\textbf{{V}}_{l}$
3 Hyper Parameters: $H$, number of heads; $B$, batch size
4 Hyper Parameters: $d_{k}$, embedding size; $d_{h}=H\times d_{k}$
5 $\textbf{{X}}=\textbf{{X}}.view(B,N,H,d_{k}).transpose(1,2)$
6 $\textbf{{V}}=\textbf{{V}}_{l}^{T}\times\textbf{{X}}$  # $shape=(B,H,d_{l},d_{k})$
7 $\textbf{{A}}=Softmax(\textbf{{A}}_{l},dim=-1)$  # $shape=(B,1,N,d_{l})$
@@
9 $\textbf{{Out}}=\textbf{{Out}}.transpose(1,2).view(B,N,d_{h})$
10 $\textbf{{Out}}=\textbf{{W}}^{O}(\textbf{{Out}})$
11 Output: Out  # $shape=(B,N,d_{h})$
Comparing equations (12) and (11), it is evident that the computational complexity of the vanilla self-attention module and the proposed UAM is $O\left(d_{k}\times N^{2}\right)$ and $O\left(d_{k}\times d_{l}\times N\right)$, respectively. Here, note that we omit the Q and K projections in the vanilla Transformer which results in further parameter savings. In fact, as observed in the subsequent experiments, a small value of $d_{l}$ is sufficient to achieve a good bit error rate. Therefore, the proposed UAM is more suitable for large-scale inputs and is more accommodating to long channel codes compared to vanilla self-attention module. The proposed multi-head unified-attention and corresponding Python-style pseudocode are illustrated in Figure 5 and Algorithm 1, respectively. As depicted in Figure 5, thanks to the sparsity of the attention scores, multiple heads are able to share the same attention $\textbf{{A}}_{l}$, resulting in a more cohesive decoding architecture and further reducing the complexity of the neural network. This approach contrasts with vanilla self-attention mechanisms, in which each head independently computes its scores using separate learned weights.
3.2 Sparse Masked Attention
To detect and correct errors, a decoder needs to examine and contrast the received codeword using parity check equations, where a non-zero syndrome signals the presence of channel errors. In the Transformer architecture, attention is sparse as each parity bit does not correlate with every information bit. To address this, we propose enhancing the attention model with insights from human intelligence, which can clarify the dependencies between information and parity bits. By incorporating these dependencies, the model can more rapidly and accurately grasp the underlying principles of channel decoding, leading to enhanced performance.
Specifically, for any linear block code defined by a parity-check matrix $\textbf{H}\in\mathbb{F}_{2}^{(n-k)\times n}$, we define a mapping function $M(\textbf{H}):\{0,1\}^{(n-k)\times n}\to\{-\infty,0\}^{N\times d_{l}}$. This function operates on H and produces a mask that is applied to the attention memory $\textbf{{A}}_{l}$, such that
$UniAttn\left(\boldsymbol{A}_{l},\boldsymbol{V}_{l}\right)=Softmax\left(\boldsymbol{A}_{l}+M(\textbf{H})\right)\times\left(\boldsymbol{V}_{l}^{T}\boldsymbol{X}\right)$ (13)
Let $\textbf{{A}}_{l}$ be a matrix in $\mathbb{R}^{N\times d_{l}}$, where $N$ is defined as $2n-k$ and $d_{l}$ as $n-k$. We naturally proceed to define the density of H as
$H_{d}=\frac{sum({\textbf{H}})}{(n-k)\times n}$ (14)
We then construct the extended parity-check matrix $\overline{\textbf{H}}$ as follows
$\overline{\textbf{H}}=\begin{bmatrix}{\textbf{H}}^{T}\\ {\textbf{{J}}}_{n-k}\end{bmatrix}$ (15)
where ${\textbf{{J}}}_{n-k}$ is a randomly generated sparse matrix of size $(n-k)\times(n-k)$ with a density of $H_{d}$. The generation of the attention mask, therefore, becomes a direct process and is formulated as $M(\overline{\textbf{H}})$.
@@ 3.2 Sparse Masked Attention
For enhanced clarity, Figure 6 offers a visual depiction of matrices $\textbf{H}^{T}$, $\overline{\textbf{H}}$ and the Tanner graph, utilizing the Hamming (7,4) code as an illustrative example. Here, matrix H is defined as
@@
1 Input: H  # $shape=(n-k,n)$
2 $H_{d}=\frac{sum(\textbf{H})}{(n-k)\times n}$
3 $\overline{\textbf{H}}=zeros(2n-k,n-k)$
4 $\overline{\textbf{H}}[0:n,0:n-k]=\textbf{H}.transpose(0,1)$
5 $\overline{\textbf{H}}[n:2n-k,0:n-k]=rand(n-k,n-k)<H_{d}$
@@ 6 Output: ()(¬$(-\infty)\cdot(\neg\mathrm{~{}\overline{\textbf{H}}})$
This illustration demonstrates that the application of the mask results in sparser attention, a phenomenon particularly pronounced in low-density parity-check codes. Most importantly, the complexity of the UAM is now reduced to the density of the code $O(H_{d}\times d_{k}\times d_{l}\times N)$. Our experimental results indicate that the use of this mask has successfully reduced the computational complexity of the attention mechanism by an average of $83\%$. Furthermore, we provide a summary of the Python-style sparse mask construction within Algorithm 2.
3.3 Architecture and Training
The architecture of the proposed unified error correction code Transformer is illustrated in Figure 7. The primary innovation in our model involves substituting the vanilla attention mechanism with a multi-head unified-attention module. This modification not only decreases computational complexity but also empowers the model to discern distinctive features among diverse codewords, thereby facilitating the establishment of a unified decoding architecture. We incorporate advanced pre-training strategies to optimize the model’s error correction performance. Furthermore, we have designed standardization units to standardize the lengths of codewords and syndromes across diverse datasets, ensuring the model’s compatibility with various code types, lengths, and rates.
@@ 3.3 Architecture and Training
We employ binary cross-entropy as the loss function with the objective of learning to predict the multiplicative noise z. The corresponding target binary multiplicative noise is denoted as $=bin(sign(\textbf{{z}}))$ z ) ). Therefore, the loss calculation for a received individual codeword y is given by
$loss=-\sum_{i=1}^{n}{\tilde{z}}_{i}log(f_{\theta}(y))+(1-{\tilde{z}}_{i})log(1-f_{\theta}(y))$ (17)
@@
where $\textbf{0}_{c}$ and $\textbf{0}_{s}$ represent zero vectors with sizes $N_{max}-size(\textbf{{y}})$ and $S_{max}-size(s(\textbf{{y}}))$, respectively. Through this unit, the Transformer neural decoder can accommodate any codeword within the maximum length, allowing for mixed training in a unified manner.
In the training phase, we conducted 1000 epochs, each consisting of 1024 minibatches, which in turn contained 512 samples each. Samples for each minibatch were randomly selected from a pool of systematic Polar[28], LDPC, and BCH codewords. The learning rate was initialized at $10^{-3}$, and a cosine decay scheduler was applied without warmup [29], gradually reducing it to $10^{-6}$ by the end of the training process. We employ the Adam optimizer to dynamically adjust the learning rate, leveraging its efficacy in managing diverse data types and model architectures [30]. The AWGN noise introduced to the links had its signal-to-noise ratio (SNR) randomly chosen between 4 and 6 for each batch. In the pre-training phase, we conducted 1000 epochs while maintaining consistency with all other parameters from the training phase. The word embeddings acquired during the pre-training stage served as the initial values for the subsequent training phase. The training and experiments were performed on NVIDIA GeForce RTX 4090 24GB GPUs.
@@ 4 Experiments
We conducted experiments on linear block codes, specifically Polar, LDPC, and BCH, to assess the effectiveness of the proposed architecture. The parity-check matrices utilized in the paper were sourced from [31]. We benchmark our approach against the most recent state-of-the-art performance reported in [22], where each codeword parameter was trained individually and also contrast it with the legacy successive cancellation list decoding algorithm for Polar codes [32], as well as the normalized min-sum decoding algorithm for LDPC codes [10] [33]. The results are presented in the form of bit error rate (BER) and block error rate (BLER) at various normalized SNR (i.e. $E_{b}/N_{0}$) values, with at least $10^{6}$ random codewords tested for each. The terms CBER and CBLER are hereby defined to denote the comprehensive bit error rate and block error rate, respectively, for the entirety of codewords utilized, spanning Polar, LDPC, and BCH codes.
@@
The proposed unified error correction code Transformer is defined by the following hyperparameters: the number of Transformer encoder layers $L$, the dimensions of word embeddings $d_{k}$, and the dimensions $d_{l}$ of the trainable memory $\textbf{{A}}_{l}$ and $\textbf{{V}}_{l}$ along with the number of unified attention heads $H$. We present the performance of our framework for four different architectures with $L\in\{3,6\},d_{k}\in\{16,32\}$, while all other hyperparameters are set to their default values: $d_{l}=64,H=8$. The hyperparameters in [22] are set as follows: $L=6,d_{k}=32,H=8$. Additionally, for Polar SCL decoding, the list size is set to 8, and for LDPC NMS decoding, the maximum number of iterations is set to 15 with a normalized factor of 0.875.
Table 1 presents the simulation results, specifically showing the negative logarithm base 10 of the BER at $E_{b}/N_{0}$ values of 4, 5, and 6 dB. It is noted that higher values in this context are indicative of better performance. The best performance for each SNR is highlighted in bold and italics. We provide additional BER plots for Polar and LDPC codes in Figure 8 and Figure 9, respectively. Experimental results indicate that our approach achieves comparable performance to models trained individually as reported in [22], and outperforms both the SCL algorithm for Polar codes and the NMS algorithm for LDPC codes. Our proposed unified code-agnostic Transformer effectively decodes a wide range of code types, lengths, and rates. Especially for short LDPC codes, traditional NMS algorithms are constrained by a smaller girth, leading to poor performance because LLR information cannot propagate effectively during the iteration process. In contrast, our approach sidesteps this issue by eliminating iteration, thereby offering a practical solution for short LDPC code applications, such as those for future 6G mobile communication control channel coding schemes. Furthermore, the performance gains observed with smaller dimensions $d_{l}=64\ll N=192$ confirm that the proposed decoding architecture substantially reduces computational complexity.
Table 1: A comparison of the negative logarithm base 10 of the BER
Figure 8: BER for Polar codes under $L=6,d_{k}=32$.
Figure 9: BER for LDPC codes under $L=6,d_{k}=32$.

### v2

@@ 3 Proposed Unified Error Correction Code Transformer
3.1 Shared Unified Attention
The vanilla self-attention module can be viewed as establishing correlations among different input features within a data sample while neglecting the intrinsic relationships between different samples. This limitation hinders its ability to learn the distinctiveness of features between various code types in channel decoding tasks. Furthermore, research in [30] suggests that the input of the softmax function is probabilistically sparse, indicating that only a few weights dominate. This implies that an $N\times N$ self-attention matrix is not strictly necessary. This sparsity suggests that a low-rank approximation could effectively capture the dominant correlations without sacrificing performance.
Figure 5: Rank distribution of the $\textbf{{Q}}\cdot\textbf{{K}}^{T}$ matrix in the ECCT model across 6 encoder layers for various code types.
To validate this, we trained an ECCT model (with parameters $L=6$, $H=8$, $d_{k}=32$, $d_{f}=1024$) on a variety of linear block codes, including Polar(32, 16), Polar(64, 32), Polar(128, 86), LDPC(49, 24), LDPC(121, 60), LDPC(121, 80), BCH(31, 16), BCH(63, 36), and BCH(127, 120). We analyzed the rank distribution of the $\textbf{{Q}}\cdot\textbf{{K}}^{T}$ matrix across 1000 batches for the 6-layer Transformer encoder. As shown in Figure 5, the rank ranges from 11 to 32, consistently indicating a low-rank structure across all layers and code types.
Based on the aforementioned considerations, we have developed a novel, low-complexity Unified Attention Module (UAM) to tackle channel decoding tasks involving multiple code types, lengths, and rates. We introduced learnable memory components, $\textbf{{A}}_{l}$ (attention) and $\textbf{{V}}_{l}$ (values), which play the role of learning features from the entire training dataset. The mathematical expression is given by:
$\text{UniAttn}\left({{\textbf{{A}}_{l}},{\textbf{{V}}_{l}}}\right)=\text{Softmax}\left({{\textbf{{A}}_{l}}}\right)\cdot\left({{\textbf{{V}}_{l}^{T}}\textbf{{X}}}\right)$ (9)
where $\textbf{{X}}\in\mathbb{R}^{N\times d_{k}}$ is the input of attention layer, $\textbf{{A}}_{l},\textbf{{V}}_{l}\in\mathbb{R}^{N\times d_{l}}$ are trainable parameters, and $d_{l}$ is an adjustable hyperparameter controlling the low-rank approximation.
In order to gain a better understanding of the proposed UAM, we rewrite equation (2) as follows:
$\text{Attn}\left(\textbf{{W}}^{\textit{Q}},\textbf{{W}}^{\textit{K}},\textbf{{W}}^{\textit{V}}\right)=\text{Softmax}\left(\frac{\left(\textbf{{XW}}^{\textit{Q}}\right)\cdot{\left(\textbf{{XW}}^{\textit{K}}\right)}^{T}}{\sqrt{d_{k}}}\right)\cdot\left(\textbf{{XW}}^{\textit{V}}\right)$ (10)
@@ 2 Trainable Parameters: $\textbf{{A}}_{l}$, $\textbf{{V}}_{l}$
3 Hyper Parameters: number of heads $H$; batch size $B$
4 Hyper Parameters: embedding size $d_{k}$; $d_{h}=H\times d_{k}$
5 $\textbf{{X}}=\textbf{{X}}.\text{view}(B,N,H,d_{k}).\text{transpose}(1,2)$
6 $\textbf{{A}}=\text{Softmax}(\textbf{{A}}_{l},\text{dim}=-1)$  # $\text{shape}=(B,1,N,d_{l})$
7 $_{o}=\textbf{{A}}\cdot\left(\textbf{{V}}_{l}^{T}\cdot\textbf{{X}}\right)$ X )  # $\text{shape}=(B,H,N,d_{k})$
8 $\textbf{{X}}_{o}=\textbf{{X}}_{o}.\text{transpose}(1,2).\text{view}(B,N,d_{h})$
9 $\textbf{{X}}_{o}=\textbf{{W}}^{O}(\textbf{{X}}_{o})$
10 Output: $\textbf{{X}}_{o}$  # $\text{shape}=(B,N,d_{h})$
By comparing equations (9) and (10), the computational complexity of the vanilla self-attention module is $O(N^{2}\times d_{k}+N\times d_{k}^{2})$, whereas the proposed UAM reduces it to $O(N\times d_{l}\times d_{k})$. Importantly, the omission of the Q and K projections in the vanilla Transformer leads to significant parameter savings. As demonstrated in subsequent experiments, a $d_{l}$ value comparable to $n-k$ is sufficient to achieve a favorable bit error rate. This is supported by the fact that the maximum rank of a parity-check matrix $\textbf{H}\in\{0,1\}^{(n-k)\times n}$ is $n-k$, providing a theoretical basis for the observed performance. It is also important to note that both ECCT and FECCT are based on the vanilla Transformer architecture, differing only in their self-attention mask designs.
The proposed multi-head unified-attention and corresponding Python-style pseudocode are illustrated in Figure 6 and Algorithm 1, respectively. As depicted in Figure 6, thanks to the sparsity of the attention scores, multiple heads are able to share the same attention $\textbf{{A}}_{l}$, resulting in a more cohesive decoding architecture and further reducing the complexity of the neural network. This approach contrasts with vanilla self-attention mechanisms, in which each head independently computes its scores using separate learned weights.
3.2 Sparse Masked Unified Attention
Error detection and correction in channel decoding rely on analyzing received codewords against parity-check equations, where a non-zero syndrome indicates channel errors. In Transformer architectures, vanilla self-attention is inherently dense, correlating all bits indiscriminately, despite parity bits not interacting with every information bit. To address this, we propose a sparse masked unified attention mechanism that incorporates code-specific dependencies, inspired by human intelligence in discerning bit relationships. This approach accelerates the model’s understanding of decoding principles, enhancing performance.
Specifically, for any linear block code defined by a parity-check matrix $\textbf{H}\in\mathbb{F}_{2}^{(n-k)\times n}$, we define a mapping $\text{M}(\textbf{H}):\{0,1\}^{(n-k)\times n}\to\{-\infty,0\}^{N\times d_{l}}$ that generates a mask for the attention memory $\textbf{{A}}_{l}$. The UAM is formulated as:
$\text{UniAttn}\left(\boldsymbol{A}_{l},\boldsymbol{V}_{l}\right)=\text{Softmax}\left(\boldsymbol{A}_{l}+\text{M}(\textbf{H})\right)\cdot\left(\boldsymbol{V}_{l}^{T}\boldsymbol{X}\right)$ (11)
@@ 3.2 Sparse Masked Unified Attention
$\left[\overline{\textbf{{y}}},\text{s}\left(\textbf{{y}}\right)\right]\cdot\overline{\textbf{H}}=\textbf{0}$ (12)
Inspired by this, we explore a mask structure leveraging the parity-check constraints. Let $\textbf{{A}}_{l}$ be a matrix in $\mathbb{R}^{N\times d_{l}}$, where $N$ is defined as $2n-k$ and $d_{l}$ as $n-k$. We then construct the extended parity-check matrix $\overline{\textbf{H}}$ as follows:
$\overline{\textbf{H}}=\begin{bmatrix}{\textbf{H}}^{T}\\ {\textbf{{I}}}_{n-k}\end{bmatrix}$ (13)
where ${\textbf{{I}}}_{n-k}\in\mathbb{R}^{(n-k)\times(n-k)}$ is an identity matrix with ones on the diagonal and zeros elsewhere. The generation of the attention mask is thus a direct process, formulated as $\text{M}(\overline{\textbf{H}})$.
We naturally proceed to define the density of $\overline{\textbf{H}}$ as:
$H_{d}=\frac{\text{sum}(\overline{\textbf{H}})}{(2n-k)\times(n-k)}$ (14)
@@
For enhanced clarity, Figure 7 offers a visual depiction of matrices $\textbf{H}^{T}$, $\overline{\textbf{H}}$ and the Tanner graph, utilizing the Hamming (7,4) code as an illustrative example. Here, matrix H is defined as:
@@
1 Input: H  # $\text{shape}=(n-k,n)$
2 $\overline{\textbf{H}}=\text{zeros}(2n-k,n-k)$
3 $\overline{\textbf{H}}[0:n,0:n-k]=\textbf{H}.\text{transpose}(0,1)$
4 $\overline{\textbf{H}}[n:2n-k,0:n-k]=\text{eye}(n-k)$
@@ 5 Output: ()(¬$(-\infty)\cdot(\neg\mathrm{~{}\overline{\textbf{H}}})$
This illustration demonstrates that the application of the mask results in sparser attention, a phenomenon particularly pronounced in low-density parity-check codes. Most importantly, the complexity of the UAM is now reduced to the density of the code $O(N\times d_{l}\times d_{k}\times H_{d})$. Our experimental results indicate that the use of this mask has successfully reduced the computational complexity of the attention mechanism by an average of 86%. Furthermore, we provide a summary of the Python-style sparse mask construction within Algorithm 2.
3.3 Architecture and Training Methodology
The architecture of the proposed unified error correction code Transformer is illustrated in Figure 8. The primary innovation in our model involves substituting the vanilla attention mechanism with a multi-head shared unified-attention module. This modification not only decreases computational complexity but also empowers the model to discern distinctive features among diverse codewords, thereby facilitating the establishment of a unified decoding architecture. We employ fine-tuning techniques to enhance the model’s generalization to unseen codewords. Furthermore, we have designed standardization units to standardize the lengths of codewords and syndromes across diverse datasets, ensuring the model’s compatibility with various code types, lengths, and rates.
@@ 3.3 Architecture and Training Methodology
We employ binary cross-entropy as the loss function with the objective of learning to predict the multiplicative noise z. The corresponding target binary multiplicative noise is denoted as z~=bin(sign($\textbf{{z}}))$ = bin ( sign ( z ) ). Therefore, the loss calculation for a received individual codeword y is given by:
$\text{loss}=-\sum_{i=1}^{n}{\tilde{z}}_{i}\text{log}(f_{\theta}(y))+(1-{\tilde{z}}_{i})\text{log}(1-f_{\theta}(y))$ (16)
@@
where $\textbf{0}_{c}$ and $\textbf{0}_{s}$ represent zero vectors with sizes $N_{\text{max}}-\text{size}(\textbf{{y}})$ and $S_{\text{max}}-\text{size}(s(\textbf{{y}}))$, respectively. Through this unit, the Transformer neural decoder can accommodate any codeword within the maximum length, allowing for mixed training in a unified manner.
In the training phase, we conducted 1000 epochs, each consisting of 1000 minibatches, which in turn contained 512 samples each. Samples for each minibatch were randomly selected from a pool of systematic Polar[31], LDPC, and BCH codewords. The learning rate was initialized at $10^{-3}$, and a cosine decay scheduler was applied without warmup [32], gradually reducing it to $10^{-6}$ by the end of the training process. We employ the Adam optimizer to dynamically adjust the learning rate, leveraging its efficacy in managing diverse data types and model architectures [33]. The AWGN noise introduced to the links had its Signal-to-Noise Ratio (SNR) randomly chosen between 3 and 7 for each batch. All experiments were conducted on NVIDIA GeForce RTX 4090 24GB GPUs.
@@ 4 Experiments
We conducted experiments on linear block codes, specifically Polar, LDPC, and BCH, to assess the effectiveness of the proposed architecture. The parity-check matrices utilized in the paper were sourced from [34]. We benchmark our approach against the most recent state-of-the-art performance from [24], which trains each codeword parameter independently, and the hybrid-trained FECCT [25]. Additionally, we compare it to legacy BP-based decoders, including unlearned BP [35], fully supervised Hyper BP [36], and Autoregressive BP (AR BP) [37]. The results are presented in the form of bit error rate (BER) and block error rate (BLER) at various normalized SNR (i.e. $E_{b}/N_{0}$) values, with at least $10^{5}$ random codewords tested per code type. We define aggregate BER (ABER) and aggregate BLER (ABLER) as the overall error rates spanning all evaluated codewords.
In addition, to enable direct comparison with ECCT and FECCT, whose simulation results were directly obtained from their respective papers [24, 25], we fine-tuned the pre-trained UECCT models on Consultative Committee for Space Data Systems (CCSDS) [38] and MacKay [39] codewords leveraging the hybrid-trained Polar, LDPC, and BCH models. Given GPU memory constraints and the fact that the performance of long codes has been thoroughly explored [40], while the error correction of short codes remains far from the Shannon limit [41], this work primarily focuses on training and validating medium-to-short codes.
@@
The proposed unified error correction code Transformer is defined by the following hyperparameters: the number of Transformer encoder layers $L$, the number of unified attention heads $H$, the dimensions of word embeddings $d_{k}$, the dimensions $d_{l}$ of the trainable memory $\textbf{{A}}_{l}$ and $\textbf{{V}}_{l}$, and the dimension $d_{f}$ of the FFN layer. We present the performance of our framework with hyperparameters are set to $L=6$, $H=8$, $d_{k}=64$, $d_{l}=64$, and $d_{f}=4\cdot H\cdot d_{k}$. The value of $d_{l}=64$ corresponds to the maximum syndrome length among the trained codewords.
Figure 9: BER comparison of ECCT, FECCT, and proposed UECCT for Polar codes.
Figure 10: BER comparison of ECCT, FECCT, and proposed UECCT for BCH codes.
Table 1: A comparison of the negative natural logarithm of the BER
Table 1 presents the simulation results, specifically showing the negative natural logarithm of the BER at $E_{b}/N_{0}$ values of 4, 5, and 6 dB. It is noted that higher values in this context indicate better performance, with the best results for each SNR highlighted in bold. Additional BER plots for Polar and BCH codes are provided in Figure 9 and Figure 10, respectively. Simulation results demonstrate that our proposed UECCT outperforms both the individually trained ECCT [24] and the hybrid-trained FECCT [25] across the majority of codewords, validating the effectiveness of the proposed architecture. Notably, UECCT even surpasses FECCT’s performance with an embedding length of 128 when operating at a more compact embedding length of 64, further highlighting the efficiency and power of our design. The superior performance stems from its ability to leverage a shared attention mechanism across heads, as detailed in Section 3, combined with sparsity-aware masking that reduces computational complexity while preserving decoding accuracy.